A car rental agency charges $150 per week plus $.20 per mile to rent a car. How many miles can you travel for $250?
Answer by checkley77(12569)

150+.20x=250
.20x=250-150
.20x=100
x=100/.20
x=500 miles ans.
Proof:
150=.20*500=250
150+100=250
250=250